OpenStudy (anonymous):

Solve: log(6) (x^2-5x) = 1

OpenStudy (anonymous):

that's log base 6 just to clarify

OpenStudy (anonymous):

\[\log_b(\xi)=1\iff \xi =b\]

OpenStudy (anonymous):

so \[\log_6(\xi)=1\iff \xi =6\]

OpenStudy (anonymous):

ooohhhh I know what your saayin now

OpenStudy (anonymous):

ok and then what after that

OpenStudy (anonymous):

right you only job is to solve \[x^2-5x=6\]

OpenStudy (anonymous):

wait now that doesn't make sense

OpenStudy (anonymous):

I thought they have to be same base?

OpenStudy (anonymous):

all I did was switch those and then I could just dropo the log?

OpenStudy (anonymous):

i will let you figure out why if \[\log_6(x^2-5x)=1\] then \[x^2-5x=6^1\]

OpenStudy (anonymous):

heres what ihave after I switched them